OpenStudy (anonymous):

I need the x intercept and i know it's (0,0) because my factoring is no solution, but how do i prove that?

OpenStudy (anonymous):

What do you mean your factoring is no solution?

OpenStudy (anonymous):

\[y=x ^{2}\sqrt{9-x ^{2}}\]

OpenStudy (anonymous):

or at least i think it's no solution. i know it's (0,0) b/c of the calculator, but my work isn't proving it

OpenStudy (anonymous):

Ok,so the x intercept being zero means y = 0. \[0=x^2 \sqrt{9-x^2}\]\[\implies \begin{array}{ccc} x^2 = 0 &\text{or} & 9-x^2\end{array} = 0\]

OpenStudy (anonymous):

So either x = 0 or x = 3

OpenStudy (anonymous):

Both of those are x intercepts

OpenStudy (anonymous):

shoudnt it be \[\pm3\]

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

okay just checkin.